# RPNT: Robust Pre-trained Neural Transformer - A Pathway for Generalized Motor Decoding

### v1

@@ Appendix E Data Pre-processing Pipeline
We designed a custom preprocessing pipeline to build the train dataset (for pretraining) and test dataset (for downstream evaluation) via torchbrain for the public data benchmark (Perich et al., 2018). In particular, the SpikeBinner in the pipeline converts spike timestamps to binned spike counts for transformer input, each data is 1s with 20ms bins without overlapping windows, leading to (winodws, 50, units). For the pretraining dataset, we filter the data from $intervals.start$ to $interval.end$ (thus not trial-aligned). For the downstream evaluation dataset, we use the data from $intervals.gocuetime$ and $interval.stoptime$ for CO tasks, and again use data from $intervals.start$ to $interval.end$ for RT tasks. We build the data via non-overlapping sequential bins per session, and there is no random shuffling of the datasets. Last We use the same temporal order of $70\%$, $10\%$, $20\%$ for train/val/test data split as in POSSM (Ryoo et al., 2025).
To batch the data for an efficient transformer training, we propose a simple strategy for automatic handling of various numbers of neuron units in each session. Specifically, RPNE randomly resamples the neuron activity to the maximum number of neurons. For example, assume for a single session, we have the data size with the shape (T, 48), where 48 is the number of neurons, and we require it to be consistent with (T,100) for the batch process. Thus, we do random neuron sampling of another 52 neural activities to match (T,100), e.g., (T,48) data is replicated and replicated again to repeat for the first four neurons (T, 4) to compose (T,52). Such a strategy would also apply to another session with more neurons than 100, then 100 neurons will be randomly selected to exactly match the shape (T,100). In practice, we find this approach is very powerful as our proposed uniform random spatial and temporal masking is always applied during RPNT pretraining; such a simple solution leads to robust model pretraining.
Appendix F Formulation of MRoPE
@@
For simplicity, we use 3D-RoPE as a concrete example of the general MRoPE illustration. The 4D-RoPE can be found in section F.4. For neural recordings from $S$ sites over $T$ timesteps, we require positional encodings $\mathbf{PE}(s,t,d)\in\mathbb{R}^{d_{model}}$ where $d$ is the model embedding dimension. A good positional encoding is preferred to satisfy two key properties: (1) site specific: $\mathbf{PE}(s_{i},t,d)\neq\mathbf{PE}(s_{j},t,d)$ for $i\neq j$, ensuring distinct representations for different site locations; (2) zero-shot generalization: the encoding function can generate reasonable postional embedding giving any new sites $\mathbf{s}_{new}=(x_{new},y_{new})$ (or configurations in MRoPE case).
F.1 Extension from standard RoPE to 3D-RoPE
@@
The key insight of the standard RoPE is that attention scores depend only on relative distances. Thus, combining the RoPE and session configuration (spatial coordinates $(x,y)$ for site $s$ and temporal position $t$), our goal is to build 3D-RoPE, in which the attention score can depend on relative spatial and temporal distances: $f(|x_{i}-x_{j}|,|y_{i}-y_{j}|,|t_{i}-t_{j}|)$. For position $(x,y,t)$, the complete 3D-RoPE transformation is a block-diagonal matrix:
@@
$\mathbf{R}_{m}(m)=\begin{bmatrix}\cos(m\theta^{(m}_{0})&-\sin(m\theta^{(m)}_{0})&&\\ \sin(m\theta^{(m)}_{0})&\cos(m\theta^{(m)}_{0})&&\\ &&\ddots&\\ &&&\cos(m\theta^{(m)}_{\frac{d}{6}-1})&-\sin(m\theta^{(m)}_{\frac{d}{6}-1})\\ &&&\sin(m\theta^{(m)}_{\frac{d}{6}-1})&\cos(m\theta^{(m)}_{\frac{d}{6}-1})\end{bmatrix}$ (2)
F.2 Relative Position Property Proof
@@
F.3 Other positional embedding baselines
Original RoPE:
@@ F.3 Other positional embedding baselines
Learnable PE We also implement an improved version of the traditional learnable PE (Devlin et al., 2019) so that it is 1) learnable; 2) has a simple mechanism for the configuration generalization. Thus, we employ a learned MLP that directly maps spatial-temporal coordinates to the embedding space:
@@
where $\alpha$ is a spatial scaling factor. In our experiments, we set the $\alpha=1$ for simplicity.
Sinusoidal PE Last, we also incorporate the standard sinusoidal positional embedding (Vaswani et al., 2017) in our ablation study.
F.4 4D-RoPE in the public benchmark
@@
where each entry was constructed similarly to  equation 2. Different from the Neuropixel datasets, in which the $(x,y)$ coordinates have values, the public benchmark was string-based meta information. Therefore, we used the discrete embeddings for representation. Specifically, we used i) $[0,1]$ to embed behavior types $\{\text{CO},\text{RT}\}$; ii) $[0,1,2,3]$ to embed subject $\{c,j,m,t\}$; iii) $[0,1]$ range normalization to the recording time (day/month/year). The choice of (i) is according to alphabetical order. To show that the alphabetical discrete embedding does not introduce inductive bias. We further performed another experiment by randomly changing the dataset assigned numbers to $[1,3,2,0]$ for $\{c,j,m,t\}$, and tested it on the T-RT test. Table 1 did not show an obvious performance gap in different subject embeddings.
@@
Table 1: Swipe study on subject embeddings
@@
Appendix G Lightweight decoder
Our main goal in RPNT is to get the pretrained transformer encoder, which later can be either finetuned for downstream tasks or frozen for data-driven functional connectivity visualization. The decoder is thus designed to be lightweight $\text{Decoder}_{\theta}:\mathbb{R}^{B\times T\times D}\rightarrow\mathbb{R}^{B\times T\times N}$ maps encoded representations to Poisson rate parameters through two feedforward blocks and an output projection:
$\bm{\lambda}$ $=\text{Softplus}(\text{Linear}(\mathbf{Z}_{2}))$ (10)
where each block is defined as:
$\text{FFN}_{1}(\mathbf{x})$ $=\text{Dropout}(\text{GELU}(\mathbf{x}\mathbf{W}_{1}+\mathbf{b}_{1})),\quad\mathbf{W}_{1}\in\mathbb{R}^{D\times\frac{D}{2}}$ (11)
$\text{FFN}_{2}(\mathbf{x})$ $=\text{Dropout}(\text{GELU}(\mathbf{x}\mathbf{W}_{2}+\mathbf{b}_{2})),\quad\mathbf{W}_{2}\in\mathbb{R}^{\frac{D}{2}\times\frac{D}{4}}$ (12)
$\text{Linear}(\mathbf{x})$ $=\mathbf{x}\mathbf{W}_{\text{out}}+\mathbf{b}_{\text{out}},\quad\mathbf{W}_{\text{out}}\in\mathbb{R}^{\frac{D}{4}\times N}$ (13)
The softplus activation ensures positive rate parameters suitable for Poisson likelihood modeling of neural spiking activity. Although this MLP-based decoder is lightweight compared to the sophisticated transformer decoder design in (Vaswani et al., 2017), it enables efficient training while maintaining decoding capacity.
Appendix H Pretraining Details
H.1 Comparison with the standard MAE and our objective
We compare our proposed causal masked neural modeling with the prior MAE approach in table 2.
Point Masked Autoencoding Causal Masked Neural Modeling
Information flow Bidirectional Unidirectional
Goal Interpolation Extrapolation
Application Offline (Denoising) Online (Prediction)
Table 2: Comparison between MAE and our proposed causal masked neural modeling
H.2 Cross-Site Contrastive Learning Details
@@
For each site representation $\mathbf{z}_{s_{i},t}$ from site $s_{i}$ at time $t$, we define the Positive instances and Negative instances as the same-site neural representation and the different sites neural representation. For simplicity, we first average the representation for all time $t$. The temperature $\tau$ controls the smoothness of the similarity distributions; we set $\tau=0.1$ consistently across all experiments on the private neuralpixel dataset. Thus, we contrast representations from different sites:
@@
We showed the ablation study for the contrastive loss in table 3. Our results showed that by contrastive loss, it encourages the RPNT model to learn more robust neural representations that have better downstream decoding performance.
Appendix I Compare with Baseline Models
We compared our RPNT with current decoding models as baselines. Specifically, we included the Wiener filter (Van Drongelen, 2018) and several standard machine learning models such as MLP (Glaser et al., 2020), GRU (Cho et al., 2014), S4D (Gu et al., 2021), Mamba (Gu and Dao, 2023), transformer (Vaswani et al., 2017; Devlin et al., 2019) that were trained from scratch on the downsrteam single session. We also included two baseline pretrained models: NDT2 (Ye et al., 2023) and the latest POSSM (Ryoo et al., 2025). We acknowledge the author’s contribution (Ryoo et al., 2025) for implementing and testing all the baseline models in the main paper Table 1.
Appendix J Visualization
We examined the attention-based FC map $\mathbf{C}\in\mathbb{R}^{T\times S\times S}$ from the spatial encoder for the Neuropixel dataset. Appendix figure 1 showed the time evolution (0ms, 300ms, 600ms) of FC during the reaching period. We observed that site 6 remained quite active, whereas large site numbers, e.g., S10 and later, seemed inactive. Further, it revealed a small network interaction within the large PMd-M1 region, which provided data-driven insights into neural mechanisms of motor behavior. These results may provide insight for future intervention investigation of brain sites to verify the data-driven observations.
Appendix K Sweeping analysis for RPNT model
@@ Appendix I Compare with Baseline Models
We showed the sweeping analysis result for our RPNT model on the T-RT task, including temporal transformer layers (see table 6), attention heads (see table 6), and kernel sizes in context-based attention (see table 6). Further, the standard deviation included ablation studies tables were listed (see table 8 and table 8).
@@
Sinusoidal PE 0.8260 $\pm$ 0.0894 0.6242 $\pm$ 0.0267
RoPE 0.8226 $\pm$ 0.0941 0.6484 $\pm$ 0.0074
Learnable PE 0.8305 $\pm$ 0.0917 0.6273 $\pm$ 0.0398
MRoPE 0.8515 $\pm$ 0.1071 0.6612 $\pm$ 0.0328
Table 7: Ablation study of positional encoding methods.
Method Cross-Task (T-RT) Cross-Site (B-CS)
Standard attention 0.8024 $\pm$ 0.0896 0.5024 $\pm$ 0.0388
Context-based attention 0.8515 $\pm$ 0.1071 0.6612 $\pm$ 0.0328
Table 8: Ablation study of attention mechanisms.
Appendix L Transfer datasets performance of our pretrained RPNT model
As both datasets basically describe NHPs doing similar motor tasks with recording on similar regions. We designed two experiments to evaluate our pretrained RNPT for dataset transferring generalization. First, we apply the neuropixel pretrained RPNT model on the public dataset. We observe that RPNT has limited generalization performance (see table 10). On the other hand, when we flip the transfer and apply pretrained RPNT on the public dataset transferred to the neuropixel dataset, we observe that RPNT is able to maintain good generalization performance (see table 10). It is interesting to observe that the pretraining on the public dataset contains robust neural representations that can be adapted to the neuropixel datasets. This may be partially explained by the amount of data between the public dataset (more than 100 sessions) and the neuropixel dataset (16 sessions).
Method C-CO T-CO T-RT
Transfer from neuropixel to public 0.8603 $\pm$ 0.0009 0.7585 $\pm$ 0.0344 0.6395 $\pm$ 0.0892
No transfer, only on public 0.9894 $\pm$ 0.0037 0.9626 $\pm$ 0.0059 0.9626 $\pm$ 0.0059
Table 9: Model Transfer from neuropixel to public
Method B-CS
Transfer from public to neuropixel 0.6537 $\pm$ 0.0413
No transfer, only at neuropixel 0.6612 $\pm$ 0.0328
Table 10: Model transfer from public to neuropixel
Appendix M Sweep analysis of the train split ratio in the neuropixel dataset
In table 11, we provide more results by sweeping the train split ratio from $1\%,5\%,10\%,20\%,30\%,40\%,50\%]$ and the sweeping results are shown as follows. In general, the pretrained RPNT model outperformed other baselines.
Method 1% 5% 10% 20% 30% 40% 50%
Wiener filter 0.0009 0.0344 0.0892 0.3462 0.3567 0.3597 0.3634
MLP 0.1181 0.3070 0.3704 0.4074 0.4215 0.4279 0.4314
RNN -0.0049 -0.0215 0.1083 0.5015 0.5534 0.6229 0.6295
NDT -0.0038 0.0836 0.2401 0.5272 0.5752 0.6082 0.6175
PoYo -0.0136 0.2285 0.5378 0.5944 0.6578 0.6938 0.7285
RPNT (pretrained) 0.0701 0.3735 0.5418 0.6612 0.6927 0.7170 0.7329
Table 11: Sweep analysis of the train split ratio in the neuropixel dataset
Appendix N Neuropixel Data Collection and Preprocessing
N.1 Neuralpixel Data Collection
We utilized high-density Neuropixel recordings (see appendix figure 2 for neuropixel setup and appendix table 13 for each recording information) from a NHP performing center-out reaching tasks. The dataset is comprised of 17 recordings at different sites distributed across the PMd and M1 regions across experiment sessions spanning approximately one year.
N.2 Preprocessing Pipeline
We required batch-like neural data for transformer model pretraining. However, the raw data across heterogeneous recording sites have varying trial counts, neuron populations, and recording times. Let $\mathcal{S}=\{S_{1},S_{2},\ldots,S_{K}\}$ denote a collection of $K$ recording sites, where each site $S_{i}$ contain neural spike data $\mathbf{X}_{i}\in\mathbb{R}^{C_{i}\times T_{i}\times N_{i}}$ with $C_{i}$ counts of successful trials, $t_{j}$ time points, and $N_{i}$ recording neurons after spike sorting. Our objective was to construct a unified dataset $\mathcal{D}=\{(\mathbf{X}_{k=1}^{K}),\mathbf{L}\}$ where $\mathbf{X}^{(k)}\in\mathbb{R}^{C\times T\times N}$ represented standardized neural activity for each site and $\mathbf{L}\in\mathbb{R}^{L\times 2}$ encode spatial location (coordinates).
Therefore, we implemented a systematic approach to load multi-site neuropixel recordings. Given a metadata file containing site locations $\mathbf{L}_{i}=(x_{i},y_{i})\in\mathbb{R}^{2}$ and corresponding data files, we filter sites based on data availability and quality criteria. Each site’s spike data is loaded as $\mathbf{X}_{i}\in\mathbb{R}^{C_{i}\times T_{i}\times N_{i}}$. We first extract only the 1s reaching period data and apply it with $20ms$ bin size. This process gives us uniform temporal dimension ($T=50$). Next, we standardize neuron numbers across heterogeneous sites while creating data variety through neuron subset sampling. For each train/valid/test split $s$ (based on trials) and site $i$ with shape $\mathbf{X}_{i}^{(s)}\in\mathbb{R}^{C_{i}^{(s)}\times T\times N_{i}}$, we apply the following pipeline:
Algorithm 1 Data Preprocessing Pipeline on the Neuropixel Dataset
2:Step 2: Neuron Multi-Sampling
[…]
4:if $N_{i}\geq N_{\text{total}}$ then
5:  $\bm{\nu}\leftarrow\text{sample\_without\_replacement}(N_{i},N_{\text{total}})$
6:else
7:  $\bm{\nu}\leftarrow\text{sample\_with\_replacement}(N_{i},N_{\text{total}})$
8:end if
[…]
11:Step 3: Trial Sampling for Target Matching
[…]
13:if $T_{\text{available}}\geq T_{\text{target}}^{(s)}$ then
14:  $\bm{\tau}\leftarrow\text{sample\_without\_replacement}(C_{\text{available}},C_{\text{target}}^{(s)})$
15:else
16:  $\bm{\tau}\leftarrow\text{sample\_with\_replacement}(C_{\text{available}},C_{\text{target}}^{(s)})$
17:end if
[…]
After neuron multi-sampling, we apply split-specific trial sampling to match target trial counts:
$C_{\text{target}}^{(s)}=\begin{cases}C_{\text{train}}&\text{if }s=\text{train}\\ C_{\text{min}}&\text{if }s\in\{\text{val},\text{test}\}\end{cases}$ (15)
where $T_{\text{train}}$ is the target training sample count and $T_{\text{min}}$ is the minimal validation/test count to prevent oversampling. Last, we aggregate standardized site data into consistent dimension tensors
$\mathbf{X}_{\text{sites}}^{(s)}$ $=\text{stack}([\hat{\mathbf{X}}_{1}^{(s)},\hat{\mathbf{X}}_{2}^{(s)},\ldots,\hat{\mathbf{X}}_{K}^{(s)}])\in\mathbb{R}^{K\times C_{\text{target}}^{(s)}\times T\times N}$ (16)
$\mathbf{X}^{(s)}$ $=\text{transpose}(\mathbf{X}_{\text{sites}}^{(s)},(1,0,2,3))\in\mathbb{R}^{C_{\text{target}}^{(s)}\times K\times T\times N}$ (17)
Our data processing approach ensures strict separation ($\mathbf{X}_{i}^{\text{train}}\cap\mathbf{X}_{i}^{\text{val}}\cap\mathbf{X}_{i}^{\text{test}}=\emptyset,\quad\forall i\in\{1,\ldots,K\}$) while tensors maintain consistent dimensionality, i.e., $\mathbf{X}^{(s)}\in\mathbb{R}^{B^{(s)}\times K\times C\times N},\quad\forall s\in\{\text{train},\text{val},\text{test}\}$, where $B^{(s)}=\tilde{C}^{(s)}$ represents the batch dimension for split $s$.
Raw trial counts range from 200 to 1417 per site; raw neuron counts range from 83 to 703 per site. Our preprocessing pipeline standardizes all site data, which generates 80,000 (16*5000) training samples for our RPNT pretraining and efficient validation/test data. Further, our pipeline generates the dataset being agnostic of neuron numbers and neuron orders; rather, it only depends on the target neurons $N$ that one can specify. Therefore, our model trained on this dataset can handle arbitrary numbers of neurons via the sampling strategy in our pipeline. To this end, table 13 and table 13 summarized the pretraining data statistics before and after the preprocessing.
Site ID Coordinates Trials Time bins Neurons Site labels
9940 (-1, 1) 1286 50 328 S13
10802 (3, 2) 1200 50 703 S11
10812 (-3, 2) 1200 50 511 S9
10820 (-2, 4) 1086 50 394 S3
10828 (-1, 4) 1417 50 329 S4
12269 (3, -4) 752 50 252 S16
12290 (5, 2) 763 50 367 S12
13153 (1, 5) 415 50 275 S2
13122 (2, 4) 674 50 83 S6
13239 (-1, 5) 967 50 154 S1
13256 (2, -4) 1024 50 255 S15
13272 (-1, 3) 1056 50 187 S7
14116 (Test) (3, -2) 789 50 258 S17
14139 (2, 2) 865 50 125 S10
14824 (2, 3) 200 50 354 S8
14878 (3, -3) 200 50 309 S14
14891 (0, 4) 200 50 106 S5
Table 12: Site-by-site breakdown showing data characteristics
Split Samples Shape Sampling Strategy
Training 80,000 $(5000,16,50,50)$ Full sampling
Validation 3,200 $(200,16,50,50)$ Minimal sampling
Test 3,200 $(200,16,50,50)$ Minimal sampling
Table 13: Dataset split characteristics after preprocessing. Format: (Trials, Sites, Time, Neurons).
Appendix O Hyperparameters and Computational resources
Parameter Microelectrode benchmark Neuropixel dataset

### v2

@@ 1 Micro-electrode Data Pre-processing Pipeline
We design a custom preprocessing pipeline to construct the training dataset for pretraining and the test dataset for downstream evaluation via torchbrain for the LTRCH benchmark [39, 40]. In particular, we create a SpikeBinner module to convert spike timestamps into binned spike counts for transformer input, where each sample spans 1s using 20ms bins without overlap, resulting in tensors of shape (windows, 50, units). For the pretraining dataset, we filter the data from $intervals.start$ to $intervals.end$, and thus the data is not trial-aligned. For the downstream evaluation dataset, we use data from intervals.gocuetimeintervals.go\ cue\ time to $intervals.stoptime$ for center out tasks, and again use data from $intervals.start$ to $intervals.end$ for random target tasks. We construct the data using non-overlapping sequential bins within each session; we do not randomly shuffle the dataset to respect the temporal order. Finally, we use the same temporal split of $70\%$, $10\%$, $20\%$ for train/val/test splits as in POSSM [43].
To batch the data efficiently for transformer training, we adopt the following strategy to deal with a varying number of recorded neurons across sessions. Specifically, RPNT randomly resamples the neural activity to a fixed maximum number of neurons. For example, for a session with data of shape $(T,48)$, where 48 is the number of neurons, we resample the data to match the batch shape $(T,100)$. In this case, we randomly sample an additional 52 neuron traces to reach the target size. For instance, the $(T,48)$ data may be repeated once and then supplemented with the first four neurons $(T,4)$ to obtain $(T,52)$, which is then concatenated with the original data to form $(T,100)$. The same strategy applies to sessions with more than 100 neurons, in which case 100 neurons are randomly selected to match the fixed shape $(T,100)$. In practice, this simple resampling strategy works well because stochastic uniform spatial and temporal masking is always applied during RPNT pretraining. In our experiments, we set the fixed number of neurons to $300$.
2 Neuropixel Data Preprocessing
We require batch-like neural data for transformer model pretraining. However, raw data collected across heterogeneous recording sites vary in trial count, neuron population, and recording duration. Let $\mathcal{S}=\{S_{1},S_{2},\ldots,S_{K}\}$ denote a collection of $K$ recording sites, where each site $S_{i}$ contains neural spike data $\mathbf{X}_{i}\in\mathbb{R}^{C_{i}\times T_{i}\times N_{i}}$ with $C_{i}$ successful trials, $t_{j}$ time points, and $N_{i}$ recording neurons after spike sorting. Our objective is to construct a unified dataset $\mathcal{D}=\{(\mathbf{X}_{k=1}^{K}),\mathbf{L}\}$ where $\mathbf{X}^{(k)}\in\mathbb{R}^{C\times T\times N}$ denotes standardized neural activity for site and $\mathbf{L}\in\mathbb{R}^{L\times 2}$ encodes spatial location coordinates.
Therefore, we implement a systematic pipeline to load multi site Neuropixel recordings from the NPCS benchmark. Given a metadata file containing site locations $\mathbf{L}_{i}=(x_{i},y_{i})\in\mathbb{R}^{2}$ and the corresponding data files, we filter sites based on data availability and quality criteria. Each site’s spike data is loaded as $\mathbf{X}_{i}\in\mathbb{R}^{C_{i}\times T_{i}\times N_{i}}$. We first extract the 1s reaching period data and bin the data using a $20ms$ bin size. This process yields a uniform temporal dimension ($T=50$). Next, we standardize neuron numbers across heterogeneous sites while preserving variability through neuron subset sampling. For each train/valid/test split $s$ (based on trials) and site $i$ with shape $\mathbf{X}_{i}^{(s)}\in\mathbb{R}^{C_{i}^{(s)}\times T\times N_{i}}$, we apply the pipeline in algorithm 1.
Algorithm 1 Data Preprocessing Pipeline on the Neuropixel Dataset
2:Step 2: Neuron Multi-Sampling
[…]
4:if $N_{i}\geq N_{\text{total}}$ then
5:  $\boldsymbol{\nu}\leftarrow\text{sample\_without\_replacement}(N_{i},N_{\text{total}})$
6:else
7:  $\boldsymbol{\nu}\leftarrow\text{sample\_with\_replacement}(N_{i},N_{\text{total}})$
8:end if
[…]
11:Step 3: Trial Sampling for Target Matching
[…]
13:if $T_{\text{available}}\geq T_{\text{target}}^{(s)}$ then
14:  $\boldsymbol{\tau}\leftarrow\text{sample\_without\_replacement}(C_{\text{available}},C_{\text{target}}^{(s)})$
15:else
16:  $\boldsymbol{\tau}\leftarrow\text{sample\_with\_replacement}(C_{\text{available}},C_{\text{target}}^{(s)})$
17:end if
[…]
After neuron multi-sampling, we apply a split-specific trial sampling to match target trial counts
$C_{\text{target}}^{(s)}=\begin{cases}C_{\text{train}}&\text{if }s=\text{train}\\ C_{\text{min}}&\text{if }s\in\{\text{val},\text{test}\},\end{cases}$ (1)
where $T_{\text{train}}$ is the target training sample count and $T_{\text{min}}$ is the minimal validation/test count to prevent oversampling. Finally, we aggregate the standardized site data into tensors with consistent dimensions
$\mathbf{X}_{\text{sites}}^{(s)}$ $=\text{stack}([\hat{\mathbf{X}}_{1}^{(s)},\hat{\mathbf{X}}_{2}^{(s)},\ldots,\hat{\mathbf{X}}_{K}^{(s)}])\in\mathbb{R}^{K\times C_{\text{target}}^{(s)}\times T\times N},$ (2)
$\mathbf{X}^{(s)}$ $=\text{transpose}(\mathbf{X}_{\text{sites}}^{(s)},(1,0,2,3))\in\mathbb{R}^{C_{\text{target}}^{(s)}\times K\times T\times N}.$ (3)
Our data processing ensures strict separation between between training, validation, and test sets ($\mathbf{X}_{i}^{\text{train}}\cap\mathbf{X}_{i}^{\text{val}}\cap\mathbf{X}_{i}^{\text{test}}=\emptyset,\quad\forall i\in\{1,\ldots,K\}$) while maintaining consistent tensor dimensionality, i.e., $\mathbf{X}^{(s)}\in\mathbb{R}^{B^{(s)}\times K\times C\times N},\quad\forall s\in\{\text{train},\text{val},\text{test}\}$, where $B^{(s)}=\tilde{C}^{(s)}$ denotes the batch dimension for split $s$.
Raw trial counts range from 200 to 1417 per site, and raw neuron counts range from 83 to 703 per site. Our preprocessing pipeline standardizes all site data, yielding 80,000 (16*5000) training samples for RPNT pretraining together with corresponding validation and test sets. The resulting dataset is agnostic to neuron count and neuron ordering; instead, it depends only on the target number of neurons $N$ specified by the user. As a result, RPNT trained on this dataset can accommodate arbitrary neuron counts through the sampling strategy implemented in the pipeline. To this end, Tables 2 and Table 2 summarize the pretraining data statistics before and after preprocessing.
Site ID Coordinates Trials Time bins Neurons Site labels
9940 (-1, 1) 1286 50 328 S13
10802 (3, 2) 1200 50 703 S11
10812 (-3, 2) 1200 50 511 S9
10820 (-2, 4) 1086 50 394 S3
10828 (-1, 4) 1417 50 329 S4
12269 (3, -4) 752 50 252 S16
12290 (5, 2) 763 50 367 S12
13153 (1, 5) 415 50 275 S2
13122 (2, 4) 674 50 83 S6
13239 (-1, 5) 967 50 154 S1
13256 (2, -4) 1024 50 255 S15
13272 (-1, 3) 1056 50 187 S7
14116 (Test) (3, -2) 789 50 258 S17
14139 (2, 2) 865 50 125 S10
14824 (2, 3) 200 50 354 S8
14878 (3, -3) 200 50 309 S14
14891 (0, 4) 200 50 106 S5
Table 1: Breakdown of data characteristics by site.
Split Samples Shape Sampling Strategy
Training 80,000 $(5000,16,50,300)$ Full sampling
Validation 3,200 $(200,16,50,300)$ Minimal sampling
Test 3,200 $(200,16,50,300)$ Minimal sampling
Table 2: Characteristics of dataset splits after preprocessing. Format: (Trials, Sites, Time, Neurons).
3 Formulation of MRoPE
@@
We first use 3D-RoPE as a concrete illustration of the general MRoPE formulation, and we demonstrate the extension to 4D-RoPE later. For neural recordings from $S$ sites over $T$ timesteps, we require positional encodings $\mathbf{PE}(s,t,d)\in\mathbb{R}^{d_{model}}$ where $d$ is the model embedding dimension. A desirable positional encoding should satisfy two key properties: (1) site specific, such that $\mathbf{PE}(s_{i},t,d)\neq\mathbf{PE}(s_{j},t,d)$ for $i\neq j$, ensuring distinct representations for different site locations (2) zero-shot generalization, such that the the encoding function can generate reasonable embeddings for previously unseen sites $\mathbf{s}_{new}=(x_{new},y_{new})$.
Extension from standard RoPE to 3D-RoPE
@@
The key insight of standard RoPE is that attention scores depend only on relative distances. By combining RoPE with session configuration variables, namely spatial coordinates $(x,y)$ for site $s$ and temporal position $t$, our goal is to construct 3D-RoPE such that the attention score depends on relative spatial and temporal distances $f(|x_{i}-x_{j}|,|y_{i}-y_{j}|,|t_{i}-t_{j}|)$. For position $(x,y,t)$, the complete 3D-RoPE transformation is a block-diagonal matrix
@@
$\mathbf{R}_{m}(m)=\begin{bmatrix}\cos(m\theta^{(m}_{0})&-\sin(m\theta^{(m)}_{0})&&\\ \sin(m\theta^{(m)}_{0})&\cos(m\theta^{(m)}_{0})&&\\ &&\ddots&\\ &&&\cos(m\theta^{(m)}_{\frac{d}{6}-1})&-\sin(m\theta^{(m)}_{\frac{d}{6}-1})\\ &&&\sin(m\theta^{(m)}_{\frac{d}{6}-1})&\cos(m\theta^{(m)}_{\frac{d}{6}-1})\end{bmatrix}$. (5)
Relative Position Property Proof
@@
Other positional embedding baselines
Original RoPE.
@@
Learnable PE. We implement an improved version of the learnable PE [15] so that it has a simple mechanism for the configuration generalization. Thus, we employ an MLP that directly maps spatial-temporal coordinates to the embedding space:
@@
where $\alpha$ is a spatial scaling factor. In our experiments, we set the $\alpha=1$ for simplicity. Sinusoidal PE. we also incorporate the standard sinusoidal positional embedding [55] in our ablation study.
4D-RoPE in the public benchmark
@@
where each entry was constructed similarly to (5). Different from the Neuropixel datasets, in which the $(x,y)$ coordinates have values, the public benchmark was string-based metadata. Therefore, we used the discrete embeddings for representation. Specifically, we used i) $[0,1]$ to embed behavior types $\{\text{CO},\text{RT}\}$; ii) $[0,1,2,3]$ to embed subject $\{c,j,m,t\}$; iii) $[0,1]$ range normalization to the recording time (day/month/year). The choice of (i) is according to alphabetical order. To show that the alphabetical discrete embedding does not introduce inductive bias. We further performed another experiment by randomly changing the dataset assigned numbers to $[1,3,2,0]$ for $\{c,j,m,t\}$, and tested it on the T-RT test. Table 3 did not show an obvious performance gap in different subject embeddings.
@@
Table 3: Sweep study on subject embeddings
@@
4 Pretraining Details
Cross-Site Contrastive Learning Details
@@ 4 Pretraining Details
For each site representation $\mathbf{z}_{s_{i},t}$ from site $s_{i}$ at time $t$, we define the Positive instances and Negative instances as the same-site neural representation and the different sites neural representation. For simplicity, we first average the representation for all time $t$. The temperature $\tau$ controls the smoothness of the similarity distributions; we set $\tau=0.1$ consistently across all experiments on the NPCS dataset. Thus, we contrast representations from different sites:
@@
We showed the ablation study for the contrastive loss in Table 4. Our results showed that contrastive loss further encourages the RPNT model to learn more robust neural representations, resulting in better downstream decoding performance.
5 Compare with Baseline Models
We compared our RPNT with current decoding models as baselines. Specifically, we included the Wiener filter [54] and several standard machine learning models such as MLP [22], GRU [10], S4D [24], Mamba [23], Transformer [55, 15] that were trained from scratch on the downsrteam single session. We also included two baseline pretrained models: NDT2 [63] and the latest POSSM [43]. We acknowledge the author’s contribution [43] for implementing and testing all the baseline models in the main paper Table 1.
6 Sweeping analysis for RPNT model
@@
We showed the sweeping analysis result for our RPNT model on the T-RT task, including temporal transformer layers (see table 7), attention heads (see table 7). We further showed the sweeping analysis of the kernel size (see table 7), which indicates the data-driven optimal context historical window was around 180ms.
7 Comparison of two causal masking implementations
@@
Causal masking before the convolution 0.8474 $\pm$ 0.0902 0.6617 $\pm$ 0.019
Causal masking after the convolution 0.8515 $\pm$ 0.1071 0.6612 $\pm$ 0.0328
Table 8: Comparison of two causal masking implementations.
We showed the results on the T-RT and B-CS tasks (as they are harder compared to the C-CO and T-CO tasks) using two causal masking implementations (see table 8). Our results did not show obvious performance differences in both implementations.
8 Spatial attention map visualization in NPCS dataset
We examined the attention-based FC map $\mathbf{C}\in\mathbb{R}^{T\times S\times S}$ from the spatial encoder for the Neuropixel dataset. Appendix figure 1 showed the time evolution (0ms, 300ms, 600ms) of FC during the reaching period. We observed that site 6 remained quite active, whereas large site numbers, e.g., S10 and later, seemed inactive. Further, it revealed a small network interaction within the large PMd-M1 region, which provided data-driven insights into neural mechanisms of motor behavior. These results might provide insight for future intervention investigation of brain sites to verify the data-driven observations.
9 Hyperparameters and Computational resources
We showed the detailed hyperparameter setup for the RPNT model pretraining and SFT in table 9 and table 10, respectively. Pretraining of RPNT was trained using a single A40 GPU, consuming 3GB and 9GB of GPU memory for the public dataset and neuropixel dataset, respectively, which takes around 12 hours for training. We saved the best checkpoint based on the early stopping criteria for the downstream SFT. For the LRTCH dataset, it ran 2 hours for C-CO, T-CO, and T-RT tasks. For the NPCS dataset, it ran 15 minutes for B-CS tasks. The SFT results were reported based on the last epoch.
Parameter Public dataset (LTRCH) Neuropixel dataset (NPCS)